OpenStudy (anonymous):

sin(a+b)sin(a-b)=sin^2a-sin^2b

OpenStudy (aravindg):

Open sin(a+b) and sin(a-b)

OpenStudy (anonymous):

and end up with sin^2Acos^2B-cos^2Asin^2b?

OpenStudy (aravindg):

yes

OpenStudy (anonymous):

then whats next

OpenStudy (aravindg):

Notice that our final answer should contain only sin terms and not cosines

OpenStudy (aravindg):

So change all cos x terms in terms of sin x

OpenStudy (aravindg):

you could use the identity \[\sin^2 x+\cos^2 =1\]

OpenStudy (anonymous):

there is no addition though when i expand???

OpenStudy (aravindg):

Wait a sec ..pls I think there is a simpler way out

OpenStudy (aravindg):

Gimme a minute

OpenStudy (anonymous):

k

OpenStudy (anonymous):

What is the problem in using \(cos^2(x) = 1- sin^2(x)\) there?? @AravindG

OpenStudy (aravindg):

It seems the power becomes 4 when we use that here

OpenStudy (anonymous):

yeah thats what i though

OpenStudy (anonymous):

Oh yeah..

OpenStudy (aravindg):

working on it ...

OpenStudy (anonymous):

lo0l waterineyes your method actually worked

OpenStudy (anonymous):

thanks man. If you expand it with your identities it works :)

OpenStudy (aravindg):

How ?

OpenStudy (anonymous):

They will cancel out I think.. ??

OpenStudy (anonymous):

sin^2a - sin^2a sin^2b-sin^2b+sin^2a sin^2B and they cancel out

OpenStudy (anonymous):

\[\sin^2(A)- \sin^2(A)\sin^2(B) - \sin^2(B) + \sin^2(A)\sin^2(B)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I was just doing calculations in mind, so that is why did not pick it up..:)

OpenStudy (anonymous):

thanks man

OpenStudy (aravindg):

Ya same here ...I mixed up the A and B ..Sorry

OpenStudy (anonymous):

You are most welcome.. :) Keep it up..

OpenStudy (anonymous):

Yeah I too forgot when you say it will become 4, I did not check when you said. Earlier I had that we could get to answer by using that identity.. :P

OpenStudy (raden):

use the identity : sinP sinQ = -1/2 (cos(P+Q - cos(P-Q) so, ur equation above can be like this : sin(a+b)sin(a-b) = -1/2(cos2a - cos2b) then use the identity : cos2x = 1 - 2sin^2 x so, that can simplied be sin(a+b)sin(a-b) = -1/2(cos2a - cos2b) = -1/2 (1-2sin^2 a - (1-2sin^2 b)) = -1/2 (-2sin^2 a + 2sin^2 b) = sin^2 a - 2sin^2 b